OpenStudy (anonymous):

if x is the average of k and 10,and y is the average of k and 4,what is the average of x and y,in terms of k?

OpenStudy (amriju):

x=(k+10)/2.......y=(k+4)/2....so avg of x and y...=(x+y)/2=1/2*((k+10)/2 + (k+4)/2))=(k+7)/2

OpenStudy (anonymous):

okaay thanks

OpenStudy (anonymous):

if f(x) is a linear function such that f(2)=5 and f(4)=13 then f(x)=?